# Evaluation of tungsten influx rate using line emissions from W5+ ions in EAST Tokamak

### v2

@@ 2 Experimental setup
Figure 1 shows the operating parameters of EAST plasma discharge #100300 as a function of time. The upper part (a) displays the plasma current $I_{p}$. Part (b) shows the electron temperature of the core plasma $T_{e0}$ provided by the Thomson scattering system [23], and the line-averaged electron density $n_{e}$ measured by the hydrogen cyanide (HCN) interferometer system [24]. Part (c) depicts the heating power of ECRH, LHW, and NBI. Part (d) shows the total radiation power loss measured by the AXUV (absolute extreme ultraviolet) photodiode array [25]. Finally, part (e) shows the emission intensity of the two dominant W5+ EUV spectral lines used here for determining the tungsten total flux. As shown in part (d), the radiation power loss starts to increase at time $t=5.14$ s, as a consequence of core tungsten accumulation, which causes the plasma to cool down dramatically, as seen in part (b). Then, the plasma became unstable. The spectral lines from low-ionized tungsten ions (e) appear in the cool plasma after strong interactions between the unstable plasma and the wall at $t\approx 5.2$ s.
Figure 1: Evolution of specific parameters for the EAST discharge #100300. (a) Plasma current $I_{p}$. (b) Central electron temperature, $T_{e0}$, and line-averaged electron density, $n_{e}$. (c) Heating power of ECRH, LHW, and NBI. (d) Boundary radiation power loss. (e) W5+ emission line intensities at 382.13 Å (blue), and 394.07 Å (red).
@@ 3.1 Collisional Radiative Model
where $n_{e}$ is the electronic density, $A_{ji}$ are the radiative rate coefficient for transitions from level $j$ to level $i$. $Q_{ji}$ are the electron-impact excitation rate coefficients if $i>j$, and the deexcitation rates if $i<j$. In matricial form, Eq. (1) is
@@
$\left(\begin{array}[]{c}\frac{dn_{\sigma}}{dt}\\ \frac{dn_{\rho}}{dt}\\ \cdots\\ \frac{dn_{\mu}}{dt}\\ 0\\ 0\\ \cdots\\ 0\end{array}\right)=\left(\begin{array}[]{cccccccc}C_{\sigma\sigma}&C_{\sigma\rho}&\cdots&C_{\sigma\mu}&C_{\sigma i}&C_{\sigma j}&\cdots&C_{\sigma n}\\ C_{\rho\sigma}&C_{\rho\rho}&\cdots&C_{\rho\mu}&C_{\rho i}&C_{\rho j}&\cdots&C_{\rho n}\\ \cdots&\cdots&\cdots&\cdots&\cdots&\cdots&\cdots&\cdots\\ C_{\mu\sigma}&C_{\mu\rho}&\cdots&C_{\mu\mu}&C_{\mu i}&C_{\mu j}&\cdots&C_{\mu n}\\ C_{i\sigma}&C_{i\rho}&\cdots&C_{i\mu}&C_{ii}&C_{ij}&\cdots&C_{in}\\ C_{j\sigma}&C_{j\rho}&\cdots&C_{j\mu}&C_{ji}&C_{jj}&\cdots&C_{jn}\\ \cdots&\cdots&\cdots&\cdots&\cdots&\cdots&\cdots&\cdots\\ C_{n\sigma}&C_{n\rho}&\cdots&C_{n\mu}&C_{ni}&C_{nj}&\cdots&C_{nn}\end{array}\right)\left(\begin{array}[]{c}n_{\sigma}\\ n_{\rho}\\ \cdots\\ n_{\mu}\\ n_{i}\\ n_{j}\\ \cdots\\ n_{n}\end{array}\right)\hfill$ (33)
@@
$\left(\begin{array}[]{c}C_{i\sigma}n_{\sigma}+\cdots+C_{i\mu}n_{\mu}+C_{ii}n_{i}+\cdots+C_{in}n_{n}\\ C_{j\sigma}n_{\sigma}+\cdots+C_{j\mu}n_{\mu}+C_{ji}n_{i}+\cdots+C_{jn}n_{n}\\ \cdots\\ C_{n\sigma}n_{\sigma}+\cdots+C_{n\mu}n_{\mu}+C_{ni}n_{i}+\cdots+C_{nn}n_{n}\\ \end{array}\right)=\left(\begin{array}[]{c}0\\ 0\\ \cdots\\ 0\\ \end{array}\right)\hfill$ (42)
which is equivalent to
$\left(\begin{array}[]{c}C_{ii}n_{i}+\cdots+C_{in}n_{n}\\ C_{ji}n_{i}+\cdots+C_{jn}n_{n}\\ \cdots\\ C_{ni}n_{i}+\cdots+C_{nn}n_{n}\end{array}\right)=\left(\begin{array}[]{c}-C_{i\sigma}n_{\sigma}+\cdots-C_{i\mu}n_{\mu}\\ -C_{j\sigma}n_{\sigma}+\cdots-C_{j\mu}n_{\mu}\\ \cdots\\ -C_{n\sigma}n_{\sigma}+\cdots-C_{n\mu}n_{\mu}\end{array}\right)$ (51)
The last expression can be written in a matricial form:
${\mathbf{C}_{s}}\cdot\vec{n}_{s}=\vec{C}_{\sigma}\cdot n_{\sigma}+\cdots+\vec{C}_{\mu}\cdot n_{\mu}$ (52)
where the reduced matrix for the excited levels is
${\mathbf{C}_{s}}=\left(\begin{array}[]{ccc}C_{ii}&\cdots&C_{in}\\ C_{ji}&\cdots&C_{jn}\\ \cdots&\cdots&\cdots\\ C_{ni}&\cdots&C_{nn}\\ \end{array}\right)$ (57)
and the column vectors are
$\vec{C}_{\tau}=\left(\begin{array}[]{c}-C_{i\tau}\\ -C_{j\tau}\\ \cdots\\ -C_{n\tau}\\ \end{array}\right)$ (62)
@@
$\vec{n}_{s}$ $=$ $[{\mathbf{C}_{s}}]^{-1}\cdot\vec{C}_{\sigma}\cdot n_{\sigma}+\cdots+[{\mathbf{C}_{s}}]^{-1}\cdot\vec{C}_{\mu}\cdot n_{\mu}=$ (63)
$=$ $\vec{{\cal F}}_{\sigma}\cdot n_{\sigma}+\cdots+\vec{{\cal F}}_{\mu}\cdot n_{\mu}$
in which the column vectors
$\vec{{\cal F}}_{\tau}\equiv[{\mathbf{C}_{s}}]^{-1}\cdot\vec{C}_{\tau}$ (64)
denote the effective contribution to the population of the excited levels via excitations from the metastable $\tau$. From this point, we want to follow the usual notation used in the field, denoting the $j$ component of the vector $\vec{{\cal F}}_{\tau}$ (i.e., the effective contribution to the population of $j$ due to excitations from the metastable $\tau$) as ${\cal F}_{j\tau}$.
@@ 3.3 Derivation of S/XB formula for many metastables
which, in matricial form is written as:
@@ 3.4 Calculational Methods
We performed the evaluation of the atomic structure and the collisional and radiative rates needed to solve the collisional-radiative problem, with three separate independent calculations. We used the atomic structure computational codes hullac (Hebrew University - Lawrence Livermore Atomic Code) [28, 29], the as (AutoStructure) code [30, 31, 32], and the fac (Flexible Atomic Code) [33]. Since the three codes are well-known in astrophysics and fusion communities, we only briefly summarize their main features. hullac is a fully relativistic atomic structure package of codes that provides the transition energies and the corresponding rate coefficients, among other quantities. The code implements the relativistic parameter potential method to solve the Dirac Hamiltonian. Full multiconfiguration wavefunctions are used to compute the radiative transition rates; configuration mixing and, therefore, correlation effects are included in the calculations. The Breit interaction and quantum electrodynamic (vacuum polarization and the self-energy) corrections are treated as second-order perturbations. The AutoStructure code allows calculating energy levels, oscillator strengths, excitation, photoionization cross sections, and autoionization rates, among other quantities. These can be calculated with configuration resolution (configuration average CA), term resolution (LS coupling), or level resolution (intermediate coupling IC) using semi-relativistic kappa-averaged wavefunctions. It also allows the inclusion of mixing configurations. fac employs a fully relativistic approach to solve the Dirac equation. Quantum-electrodynamic effects, mainly arising from Breit interaction, vacuum polarization, and electron self-energy, are all included in standard procedures in the code. fac has been widely employed to interpret laboratory and astrophysical spectroscopic data like the other programs mentioned.
@@ 4.1 Atomic Structure
The ion $W^{5+}$ is isoelectronic to the Thulium ($Z=69$), having a ground configuration [[[[Xe]]]] $4f^{14}5d$. This configuration has two levels separated by about 1 eV of energy. The lowest level of the ground configuration is $(4f^{14}5d)_{\frac{3}{2}}$ and the second level is $(4f^{14}5d)_{\frac{5}{2}}$. For most of the relevant electronic densities and temperatures, both of these two levels are significantly occupied and in particular, the population of the $J=\frac{5}{2}$ is higher than those of the ground level. We found two radiative transitions well separated and distinguishable in the experimental spectra, corresponding to the $(4f^{14}5f)_{\frac{5}{2}}\rightarrow(4f^{14}5d)_{\frac{3}{2}}$ transition at $\lambda_{1}=382.13$, and $(4f^{14}5f)_{\frac{7}{2}}\rightarrow(4f^{14}5d)_{\frac{5}{2}}$ transition at $\lambda_{1}=394.07$.
@@
Figure 3: EAST spectrum and theoretical collisional-radiative synthetic spectra for the $5f\rightarrow 5d$ transitions of the W5+ ion. (a) EAST Experimental spectra (#100300, at $t=5.200$). (b) HULLAC calculation. (c) AS (d) FAC.
@@ 4.2 Excitation and Radiative Rates
We also have calculated the radiative transition rate coefficients with the three codes, obtaining values that agree in a $20\%$ for the dipole-allowed transitions. The values of the radiative transitions and the electron-impact excitation rate coefficients for the four levels involved in the $5f\rightarrow 5d$ lines are listed in Table 2, for hullac, autostructure, and fac calculations.
@@ 4.5 Tungsten Influx
The absolute calibration of the spectral line intensities [19, 21, 22, 20] results in values of the $5f\rightarrow 5d$ transitions about $3\times 10^{13}$ photons per cm2 and per second at the emission peak corresponding to the tungsten burst. Assuming that the tungsten ions are sputtered from a region in which the electronic temperature is $T_{e}=60$ eV and the electronic density is $n_{e}=10^{13}$ cm-3, the S/XB ratios are of the order of 3, we can infer a total tungsten flux of the order of $10^{14}$ particles per cm2 and per second at the peak. Indeed, we show in Fig. 10 the calculated total tungsten flux from the plasma-facing surface towards the interior of the EAST device, which follows the behavior of the line intensity in time for discharge #100300, having a peak at $t=5.2$ seconds. The fluxes are calculated using three independent calculations: hullac, as, and fac. The three calculatons agree at the peak between $45\%$ in the worst comparison case (hullac vs. fac).
@@ 5 Conclusions
We have developed the equations for calculating the S/XB ratios in a general case where many metastable levels may contribute to the ionization of the ion. We do not assume that the populations of the metastable terms are in quasistatic equilibrium with the population of the ground term. This approach eliminates the dependence of our results on the populations of the long-lived states.
@@
Using these data, we solved a collisional-radiative model to obtain the effective population coefficients needed in to determine the S/XB ratios. We selected the radiative transitions $(4f^{14}5f)_{5/2}\rightarrow(4f^{14}5d)_{3/2}$ and $(4f^{14}5f)_{5/2}\rightarrow(4f^{14}5d)_{3/2}$, which are prominent in the 320-460 Å range. The process of radiative cascade from higher energy terms than the upper terms involved in the radiative transition was found to have a negligible effect on the SXB ratios for both types of transitions at low and intermediate energy densities. Therefore, by using a simple four-level collisional-radiative model consisting of the two upper levels and the two lower metastables, we can generate S/XB ratios that agree very well with the 430-level model. That is particularly useful because it allows us to derive simple analytical expressions that help us understand the principal mechanisms contributing to the ionization.
@@ 6 Acknowledgments
This work was supported by the National Magnetic Confinement Fusion Energy R & D Program of China (Grant Nos. 2022YFE03180400, 2022YFE03020004, 2019YFE030403), National Natural Science Foundation of China (Grant No. 12322512), and Chinese Academy of Sciences President’s International Fellowship Initiative (PIFI) (Grant Nos. 2024PVA0074 ). DM acknowledges partial support from CONICET by Project No. PIP11220200102421CO, and the ANPCyT by Project No. PICT- 2020-SERIE A-01931 in Argentina, and the Alliance of International Science Organizations (ANSO) Visiting Fellowship (ANSO-VF-2021-03), in China.

### v4

@@ 2 Experimental setup
Figure 1 (left part) displays the operating parameters of EAST plasma discharge #100300 as a function of time. The upper part (a) displays the plasma current $I_{p}$. Part (b) presents the electron temperature of the core plasma $T_{e0}$ provided by the electron cyclotron emission (ECE) system [23] and the line-averaged electron density $n_{e}$ measured by the hydrogen cyanide (HCN) interferometer system [24]. Part (c) depicts the heating power of ECRH, LHW, and NBI. Part (d) shows the total radiation power loss measured by the AXUV (absolute extreme ultraviolet) photodiode array [25]. Finally, part (e) presents the emission intensity of the two dominant W5+ EUV spectral lines at 382.13 Å and 394.07 Å, which are used here to determine the total flux of tungsten. These two lines are not frequently observed in general discharges. As the figure indicates, they do not appear during the initial and flat-top stages of the discharge, likely due to a significant reduction in throughput at longer wavelengths in the current grazing-incidence EUV spectrometer. However, at $t\approx 5.2$ s, high-intensity W5+ lines are detected. This is attributed to strong plasma-wall interactions (PWI) that ultimately lead to plasma termination. Therefore, we focus on the plasma behaviour at this final stage of the discharge, during the time interval $t=5.10-5.25$, as depicted in the right part of Figure 1. As shown in part (d), the radiation power loss begins to increase at time $t=5.14$ s due to core tungsten accumulation. This accumulation causes the plasma to cool down dramatically, as illustrated in part (b). When the electron temperature is relatively low, for instance, when $T_{e0}\leq 1.0$ keV, the emission volume of the W5+ lines increases, and meanwhile, the tungsten source increases mainly due to this strong PWI, then high-intensity spectral lines of W5+ at 382.13 Å and 394.07 Å could be observed, as shown in part (e).
Figure 1: Evolution of specific parameters for the EAST discharge #100300. (a) Plasma current $I_{p}$. (b) Central electron temperature, $T_{e0}$, and line-averaged electron density, $n_{e}$. (c) Heating power of ECRH, LHW, and NBI. (d) Boundary radiation power loss. (e) W5+ emission line intensities at 382.13 Å (blue), and 394.07 Å (red). Right Part: Zoom at $t=5.10-5.25$ s.
@@ 3.1 Collisional Radiative Model
where $n_{e}$ is the electronic density, $A_{ji}$ are the radiative rate coefficient for transitions from level $j$ to level $i$. $Q_{ji}$ are the electron-impact excitation rate coefficients if $i>j$, and the deexcitation rates if $i<j$. In matrix form, Eq. (1) is
@@
($\ \frac{d\vec{n}_{\tau}}{dt}\\ {}\hfil\\ \hline\cr{}\hfil\\ {}\hfil\\ \frac{d\vec{n}_{s}}{dt}=0\\ {}\hfil\\ {}\hfil\\ \end{array}\right)=\left(\begin{array}[]{ll|ccccc}&&&&&&\\ &&&&&&\\ &&&&&&\\ \hline\cr&&&&&&\\ &&&&&&\\ &{\mathbf{C}_{\tau}{}}&&&{\mathbf{C}_{s}}&&\\ &&&&&&\\ &&&&&&\end{array}\right)\left(\begin{array}[]{c}{}\hfil\\ \vec{n}_{\tau}\\ {}\hfil\\ \hline\cr{}\hfil\\ {}\hfil\\ \vec{n}_{s}\\ {}\hfil\\ {}\hfil\\ \end{array}\right)$ (33)
@@
00 0 $=$ ${\mathbf{C}_{\tau}}\cdot\vec{n}_{\tau}+{\mathbf{C}_{s}}\cdot\vec{n}_{s}$ (34)
@@
$\vec{n}_{s}$ $=$ $[{\mathbf{C}_{s}}]^{-1}\cdot[-{\mathbf{C}_{\tau}}]\cdot\vec{n}_{\tau}\equiv n_{e}\,{\mathbf{\cal F}}\cdot\vec{n}_{\tau}\,.$ (35)
The matrix ${\mathbf{\cal F}}$ denotes the effective contribution to the population of the excited levels via excitations from the metastables. This is the usual notation used in the field, in which ${\cal F}_{i\rho}$ represents the effective contribution to the population of $i$ due to excitations from the metastable $\rho$.
@@ 3.2 S/XB – Ionization Events per Photon
Equation (41) implicitly captures the general dependence of the S/XB coefficients on plasma parameters. As the name suggests (and will be shown below), the S/XBs are proportional to the ionization rates and inversely proportional to the excitation rates. This roughly determines their dependence on electronic temperature: while both rates increase with temperature in the relevant ranges, the effect of ionization on temperature is more pronounced. Consequently, the coefficients tend to exhibit a monotonic increase with temperature. Concerning their dependence on electron density, the S/XB coefficients appear to be independent of density at low values. At intermediate density values, the relationship becomes cumbersome, as the excitations occur through an effective rate $\mathcal{F}$ described in expression (39), which requires solving a complex collisional-radiative problem that accounts for all cascade processes.
@@ 3.3 Derivation of S/XB formula for many metastables
which, in matrix form is written as:
@@ 3.4 Calculational Methods
We carried out an evaluation of the atomic structure, as well as the collisional and radiative rates required to solve the collisional-radiative problem, using three separate independent calculations. Specifically, we utilized the atomic structure codes hullac (Hebrew University - Lawrence Livermore Atomic Code) [28, 29], the as (AutoStructure) code [30, 31, 32], and the fac (Flexible Atomic Code) [33]. Since these three codes are well-established in both astrophysical and fusion research, we will only provide a brief overview of their key features.
The hullac code is a fully relativistic atomic structure package that computes transition energies and corresponding rate coefficients, among other quantities. It employs a relativistic potential method to solve the Dirac Hamiltonian and uses full multiconfiguration wavefunctions to calculate radiative transition rates. Configuration mixing, and hence correlation effects, are included in the calculations. The Breit interaction and quantum electrodynamics corrections (vacuum polarization and self-energy) are treated as second-order perturbations.
The AutoStructure code is used to compute energy levels, oscillator strengths, excitation and photoionization cross sections, as well as autoionization rates, among other properties. These can be calculated at different levels of resolution: configuration resolution (configuration average, CA), term resolution (LS coupling), or level resolution (intermediate coupling, IC), utilizing semi-relativistic, kappa-averaged wavefunctions. The code also allows for the inclusion of configuration mixing.
Finally, the fac code uses a fully relativistic approach to solve the Dirac equation. Quantum electrodynamics effects, primarily arising from the Breit interaction, vacuum polarization, and electron self-energy, are included in the standard procedures of the code. Like the other programs mentioned, fac has been extensively used for interpreting spectroscopic data in both laboratory and astrophysical contexts.
@@ 4.1 Atomic Structure
The ion $W^{5+}$ is isoelectronic to Thulium ($Z=69$), having a ground configuration [[[[Xe]]]] $4f^{14}5d$. This configuration has two levels separated by about 1 eV of energy. The lowest level of the ground configuration is $(4f^{14}5d)_{\frac{3}{2}}$ and the second level is $(4f^{14}5d)_{\frac{5}{2}}$. For most of the relevant electronic densities and temperatures, both of these two levels are significantly occupied and in particular, the population of the $J=\frac{5}{2}$ is higher than those of the ground level. We found two radiative transitions well separated and distinguishable in the experimental spectra, corresponding to the $(4f^{14}5f)_{\frac{5}{2}}\rightarrow(4f^{14}5d)_{\frac{3}{2}}$ transition at $\lambda_{1}=382.13$, and $(4f^{14}5f)_{\frac{7}{2}}\rightarrow(4f^{14}5d)_{\frac{5}{2}}$ transition at $\lambda_{1}=394.07$.
@@
Figure 3: EAST spectrum and theoretical collisional-radiative synthetic spectra for the $5f\rightarrow 5d$ transitions of the W5+ ion. Upper graph: EAST Experimental spectra (#100300, at $t=5.200$). Lower graph: synthetic spectra calculated with HULLAC, AS, and FAC.
@@ 4.2 Excitation and Radiative Rates
We also have calculated the radiative transition rate coefficients with the three codes, obtaining values that agree $20\%$ for the dipole-allowed transitions. The values of the radiative transitions and the electron-impact excitation rate coefficients for the four levels involved in the $5f\rightarrow 5d$ lines are listed in Table 2, for hullac, autostructure, and fac calculations.
@@ 4.5 Tungsten Influx
The absolute calibration of the spectral line intensities [19, 21, 22, 20] results in values of the $5f\rightarrow 5d$ transitions about $3\times 10^{13}$ photons per cm2 per second at the emission peak corresponding to the tungsten burst. Assuming that the tungsten ions are sputtered from a region in which the electronic temperature is $T_{e}=60$ eV and the electronic density is $n_{e}=10^{13}$ cm-3, the S/XB ratios are of the order of 3, we can infer a total tungsten flux of the order of $10^{14}$ particles per cm2 per second at the peak. Indeed, we show in Fig. 10 the calculated total tungsten flux from the plasma-facing surface towards the interior of the EAST device, which follows the behavior of the line intensity in time for discharge #100300, having a peak at $t=5.2$ seconds. The fluxes are calculated using three independent calculations: hullac, as, and fac. The three calculatons agree at the peak between $45\%$ in the worst comparison case (hullac vs. fac).
@@ 5 Conclusions
In this study, we derive the relevant equations for calculating the S/XB coefficients in scenarios where, rather than assuming a single initial state, multiple metastable states with significant populations and influence on ionization processes are considered. Unlike excited states, metastable states primarily equilibrate their populations with those of neighboring ions. This allows for the expression of the relative populations of the excited states in terms of the populations of the metastable states, without the need for explicitly specifying the latter’s population values. Similarly, ionization events, which are dependent on the metastable populations, can be described in terms of the radiative emissions from the excited states.
@@
Using these data, we solved a collisional-radiative model to obtain the effective population coefficients needed in to determine the S/XB ratios. We selected the radiative transitions $(4f^{14}5f)_{5/2}\rightarrow(4f^{14}5d)_{3/2}$ and $(4f^{14}5f)_{7/2}\rightarrow(4f^{14}5d)_{5/2}$, which are prominent in the 320-460 Å range. The process of radiative cascade from energy terms higher than the upper terms involved in the radiative transition was found to have a negligible effect on the S/XB ratios for both types of transitions at low and intermediate energy densities. Therefore, by using a simple four-level collisional-radiative model consisting of the two upper levels and the two lower metastables, we can generate S/XB ratios that agree very well with the 430-level model. That is particularly useful because it allows us to derive simple analytical expressions that help us understand the principal mechanisms contributing to the ionization.
@@ 6 Acknowledgments
This work was supported by the National Magnetic Confinement Fusion Energy R & D Program of China (Grant No. 2022YFE03180400), National Natural Science Foundation of China (Grant Nos. 12322512, 12274352), and Chinese Academy of Sciences President’s International Fellowship Initiative (PIFI) (Grant No. 2024PVA0074 ). DM acknowledges partial support from CONICET by Project No. PIP11220200102421CO, and the ANPCyT by Project No. PICT- 2020-SERIE A-01931 in Argentina, and the Alliance of International Science Organizations (ANSO) Visiting Fellowship (ANSO-VF-2021-03), in China.